# Outlyingness Scores with Cluster Catch Digraphs

### v1

@@ 1 The Outlyingness Scores Based on Cluster Catch Digraphs
We have introduced several outlier detection methods based on RK-CCDs or UN-CCDs in the previous work [15],, which share a common theme: (1) Clustering Formation: Identify potential clusters and construct a single or multiple covering balls around the center, covering majority points. (2) Outlier Identification: Apply the Density-based Mutual Catch Graph (D-MCG) algorithm within each cluster to isolate points distant from the center and label them as outliers. Evaluation results demonstrated the effectiveness of these methods, with the SUN-MCCD method exhibiting superior overall performance. However, the outlyingness is not a binary property [1]. Moreover, the CCD-based methods do not provide a measure of outlyingness. An Outlyingness Score (OS) quantifies the degree of outlyingness for each observation. Incorporating OSs offers several advantages [6]: (a) OSs enhance the interpretability of outlier detection results. (b) Ranking observations by their OSs enables user-defined thresholds for outlier classification. (c) OSs contribute to a richer understanding of the data distribution and point patterns. Some established, well-known OS-based outlier detection methods include the following: (i) Isolation Forest employs multiple decision trees to compute anomaly scores [9]; (ii) Local Information Graph-based Random Walk model (LIGRW) identifies outliers by analyzing unusual patterns in random walks [17]; (iii) Local Outlier Factor (LOF) is a prototype that utilizes local reachability density to compute outlyingness [1]; (iv) Connectivity-based Outlier Factor (COF) enhances LOF by addressing limitations in detecting outliers with similar densities but differing neighbor patterns [16]; (v) LOcal Correlation Integral (LOCI) introduces a threshold-based score measuring the intensity of points within a given radius [11]; (vi) Local Outlier Probabilities (LoOP) assigns an outlier probability based on the deviates of a point from its local context [5]; (vii) Outlier Detection using In-degree Number (ODIN) identifies outliers as points with low in-degree numbers in a $k$NN graph [4]. Other OS-based methods include Angle-Based Outlier Detection (ABOD) [7], Histogram-based Outlier Score (HBOS) [3], and Feature Bagging [8].
@@ 1.1 Outbound Outlyingness Score
Given a data set $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}$ and a CCD constructed on it, the outbound neighbors of a point $x_{i}\in\mathcal{X}$, denoted as $N_{O}(x_{i})$, are those points that are covered by the covering ball centered at $x_{i}$, i.e.,
@@
Given a data set $\mathcal{X}$ and a point $x_{i}\in\mathcal{X}$, the vicinity density of $x_{i}$, denoted by $\rho_{x_{i}}$, is defined as follows,
$\rho_{x_{i}}:=\frac{|\{x_{j}|x_{j}\in B(x_{i},r_{x_{i}})\}|}{r_{x_{i}}}^{\frac{1}{d}}.$ (2)
@@
Given a data set $\mathcal{X}$ and a point $x_{i}\in\mathcal{X}$, an Outbound Outlyingness Score of $x_{i}$, denoted as $OOS(x_{i})$, is given as follows,
@@ 1.1.1 An illustration of OOS with UN-CCDs
In Figure 1(b), we construct UN-CCD to illustrate the computation of OOS. While OOS is independent of the clustering result, it is worth noting that the UN-CCD method assigns outliers $O_{1},O_{2},...,O_{6}$ to $C_{1}$, $O_{7}$ and $O_{8}$ to $C_{3}$, and $O_{9}$ to $C_{2}$. We first focus on the cluster $C_{1}$. Figure 1(b) presents the covering balls of $O_{5}$ and $O_{6}$, along with their outbound neighbors $P_{i}\ (i=1,2,3)$. $P_{i}^{\prime}s$ vicinity densities are substantially higher than those of $O_{5}$ and $O_{6}$, resulting in OOSs values of 2.47 and 10.14 for $O_{5}$ and $O_{6}$, respectively. In contrast, most of the remaining points within $C_{1}$ exhibit OOS values below 2 (Figure 1(d)). Consequently, $O_{5}$ and $O_{6}$ are clearly distinguished by their markedly higher OOS values within $C_{1}$. Similarly, $O_{9}$ receives a much higher OOS than the other points in $C_{2}$
@@
However, as a shortcoming: OOS is not effective on collective outliers, showing little robustness to the masking problem.
@@ 1.2 Inbound Outlyingness Score
Note that the larger $CI(x_{i})$, the less $x_{i}$ deviates from other points (i.e., the more likely $x_{i}$ belongs to the cluster). So, we may define the Inbound Outlyingness Score (IOS) of $x_{i}$ as the reciprocal of $CI(x_{i})$. Furthermore, to avoid zero denominator when $CI(x_{i})$ is 0 (i.e., when a point does not have any inbound neighbors), we add $\rho(x_{i})$ to $CI(x_{i})$, ranking the points without any inbound neighbors by the reciprocal of their vicinity densities.
@@ 1.2.1 An illustration of IOS with UN-CCDs
By the “three-sigma rule” [10], points with standardized scores larger than 3 can typically be deemed as outliers. However, this traditional measure is not robust to outliers, as the scores of outliers have a substantial and unbounded influence on the sample mean and the sample SD [10], which is particularly problematic in an outlier detection method. Fortunately, there are robust alternatives to mean and SD, and we employ median (Med) and the Normalized Median Absolute Deviation about the median (MADN). With this notion, a robust version of Equation (7) can be defined as follows.
@@
where $MADN(IOS(C))=Med\{IOS(C)-Med(IOS(C))\}/0.6745$. It is worth noting that the MADN of a random variable following $N(\mu,\delta)$ distribution is $\delta$ [10]. In the remaining sections, all IOS values will be standardized by default, i.e., $IOS(x_{i})=IOS^{std}(x_{i})$.
@@
Suppose the IOSs of a cluster $C$ are ranked an ascending order as follows:
@@
where the IOS of, WLOG, $x_{1}$, $x_{2}$, …, $x_{m}$ equal to $IOS_{(k)}$ ($m\leq n_{c}$). To break ties, we assign new scores to them in a linearized fashion using $\rho_{x_{1}}$, $\rho_{x_{2}}$, …, $\rho_{x_{m}}$. Without loss of generality, for any point $x_{i}$ ($i=1,...,m$), the standardized IOS of it, denoted as $\widetilde{IOS}(x_{i})$, is defined as follows,
@@
  • Effective on collective outliers and robust to masking problem.
@@
2 Monte Carlo Experiments for Outlying Scores
We evaluate the performance of all the OSs by conducting Monte Carlo experiments with diverse simulation settings. The simulation settings are elaborated in Section 5.1 the paper titled “Outlier Detection with Cluster Catch Digraphs” [15]. These settings involve different factors (e.g., dimensionality, data set sizes, cluster volumes, etc.), which vary among data sets. Moreover, there are two major types of simulation settings, one with only uniform clusters and the other with only Gaussian clusters. The goal is to evaluate four newly proposed CCD-based OSs against four existing CCD-based methods [15], seeking evidence of performance improvements.
Preliminary simulations on artificial data identified “elbow” cutoff points, serving as the thresholds for outlier detection. These thresholds, specific to each OS and dimensionality, are presented in Tables 1 and 2. Notably, the data sets with Gaussian clusters require higher thresholds than those with uniform ones, because we want to differentiate outliers from regular data points around the tails Gaussian distributions.
@@ 2 Monte Carlo Experiments for Outlying Scores
We select True Positive Rate (TPR), True Negative Rate (TNR), Balanced Accuracies (BA), and $F_{2}$-score as the performance metrics. TPR is the proportion of outlier detected, TNR is the proportion of regular points correctly labeled, BA is the mean of TPR and TNR, $F_{\beta}$-score is the weighted harmonic mean of recall (TPR) and precision, prioritizing positive observations (outliers) [13]. Since (in our opinion) recall is more important than precision in outlier detection, we set $\beta$ to 2, making recall twice as important. We emphasis on the latter two measures for enhanced accuracy since the size of outliers and regular points are highly imbalanced. Comprehensive simulation results are provided in Tables 3 to 6. For better visualization, we also provide the simulation results in four line plots (Figures 4 to 7), presenting the performance trend with varying dimensions and data sizes.
@@
Table 3: Summary of the TPRs and TNRs of all the CCD-based OSs, with the simulation settings (with uniform clusters) elaborated in Section 5.1 of [15].
Figure 4: The line plots of the TPRs and TNRs of all CCD-based OSs, under the simulation settings (with uniform clusters) elaborated in Section 5.1 of [15].
@@
Table 4: Summary of the BAs and $F_{2}$-scores of all the CCD-based OSs, with the simulation settings (with uniform clusters) elaborated in Section 5.1 of [15].
Figure 5: The line plots of the BAs and $F_{2}$-scores of all CCD-based OSs, under the simulation settings (with uniform clusters) elaborated in Section 5.1 of [15].
@@
We first look at RKCCD-OOS and UNCCD-OOS. They exhibit similar performance across different dimensions. When $d\leq 20$, both OOSs achieve high BAs and $F_{2}$-scores, often exceeding 0.95 and 0.90, respectively, when the size of the data set ($n$) is less than or equals to 200. However, as $n$ exceeds 500, the masking problem arises due to the increasing number of outliers. This issue is particularly prominent in lower dimensions where outlier intensities are high. For instance, with $d=2$ and $n=50$, UNCCD-OOS attains a BA of 0.980 and an $F_{2}$-scores of 0.932, but they drop to 0.720 and 0.475, respectively, as $n$ reaches to 1000. Conversely, in high-dimensional settings ($d\geq 50$), both OOSs achieve higher TPRs with larger data sets. For instance, when $d=100$, the TPR of RKCCD-OOS progressively increases from 0.535 to 0.994 as $n$ increases. This trend can be attributed to the fact that in high-dimensional spaces with limited data, the vicinity densities of regular observations and outliers may not differ substantially. Recall that OOS measures the relative vicinity density of a point with its outbound neighbors, the reduced performance with smaller data sets is expected. Furthermore, when $d=50$ and $n=50$, the TPR of UNCCD-OOS is 0.957, substantially higher than that of RKCCD-OOS (0.535). This is because the covering balls of UN-CCDs are typically larger than those of RK-CCD, resulting in considerable differences between the vicinity densities between regular observations and outliers, which ultimately enhances the performance of OOS.
@@
Second, we evaluate the performance of RKCCD-IOS and UNCCD-IOS. With $d\leq 20$, both IOSs demonstrate comparable performance to OOSs when the data size $n$ is small. Moreover, an improvement is observed with larger $n$ values when compared to the two OOSs. For instance, with $d=3$ and $n\leq 200$, UNCCD-IOS achieves $F_{2}$-scores comparable to those of UNCCD-OOS (0.903, 0.914 and 0.904 versus 0.954, 0.947, and 0.922). However, when $n\geq 500$, UNCCD-IOS shows substantially higher $F_{2}$-scores than those of UNCCD-OOS (0.887, 0.848 versus 0.837, 0.709). This stems from the inherent advantage of IOS, which measures the cumulative influence a point receives from other points, and the cumulative influence of outliers is typically low compared to regular observations, making IOS more robust against the masking problem, particularly in simulation settings with many outliers. For example, with $d=2$ and $n=1000$, the $F_{2}$-scores of UNCCD-OOS and UNCCD-IOS are 0.475 and 0.833, respectively, showing a huge difference. However, RKCCD-IOS does not perform well when $d\geq 50$. For example, when $d=50$ and $n=1000$, the $F_{2}$-score of RKCCD-IOS is merely 0.605, much lower than that of RKCCD-OOS, which is 0.847. A reasonable explanation is the inherent challenges of high-dimensional clustering. As $d$ increases, the average distances between outliers and clusters grow, and the covering balls constructed by RK-CCDs are not large enough with under high dimensions [15]. Consequently, as the number of outliers increases, RK-CCDs may falsely identify a set of close outliers as a valid cluster since they are not robust to the masking problem. Furthermore, IOS identifies an outlier by within-cluster comparisons, which does not mitigate this problem.
@@
Next, we compare the simulation results of the CCD-based OSs with previously proposed CCD-based outlier detection methods (which can be found in Tables 2 and 3 of [15]).
@@
Following the analysis of simulations with uniform clusters, we explore the results obtained from simulations with Gaussian clusters (see Tables 5 and 6). Similarly, we begin by assessing the performance of the four OSs, and then compare them with the previously proposed CCD-based outlier detection methods.
@@
Table 5: Summary of the TPRs and TNRs of all the CCD-based OSs, with the simulation settings (with Gaussian clusters) elaborated in Section 5.1 of [15].
Figure 6: The line plots of the TPRs and TNRs of all CCD-based OSs, under the simulation settings (with Gaussian clusters) elaborated in Section 5.1 of [15].
@@
Table 6: Summary of the BAs and $F_{2}$-scores of all the CCD-based OSs, with the simulation settings (with Gaussian clusters) elaborated in Section 5.1 of [15].
Figure 7: The line plots of the BAs and $F_{2}$-scores of all CCD-based OSs, under the simulation settings (with Gaussian clusters) elaborated in Section 5.1 of [15].
@@
Next, we consider RKCCD-IOS and UNCCD-IOS. Both IOSs achieve high $F_{2}$-scores when $d\leq 20$, because both IOSs address the masking problem raised by high-intensity outliers well. However, the performance of RKCCD-IOS drops substantially when $d\geq 50$, particularly when $n\geq 500$, For instance, when $d=100$ and $n=1000$, the $F_{2}$-score of RKCCD-IOS is 0.521. This limitation stems from the same issue encountered when using RK-CCDs for clustering, as detailed in [15].
Then, we compare the simulation results of the OSs with the U-MCCD and UN-MCCD methods. (whose performance can be found in Tables 4 and 5 of [15])
@@
Next, we compare UNCCD-OOS, UNCCD-IOS, and the UN-MCCD method, which show varying performance. The UN-MCCD method, while superior to U-MCCD, exhibits limitations because it constructs a single covering ball for each cluster, which is inadequate to cover a Gaussian cluster. In contrast, both UNCCD-OOS and UNCCD-IOS demonstrate substantial improvement. Although the $F_{2}$-score of UNCCD-OOS may be as low as 0.293 (e.g., when $d=2$ and $n=1000$) due to the masking problem under lower dimensions), and occasionally underperforms compared to the UN-MCCD method, UNCCD-IOS consistently outperforms the UN-MCCD method and achieves considerably higher $F_{2}$-scores across all the simulation settings.
@@
Based on our simulation results, we observe that RKCCD-OOS and UNCCD-OOS deliver good performance except when $d\leq 5$ and $n$ is large, where the masking problem emerges. In contract, RKCCD-IOS and UNCCD-IOS perform much better with lower dimensions but may degrade when $d$ is large because both RK-CCDs and UN-CCDs (especially RK-CCDs) may identify a set of close outliers as valid clusters in high dimensions. Compared to the U-MCCD and UN-MCCD methods, the CCD-based OSs return fewer false positives, achieving much higher $F_{2}$-scores in general. Moreover, they show substantial advantages in higher-dimensional data sets or the data sets with Gaussian clusters, making them better choices in those cases.
3 Monte Carlo Experiments Under Random Cluster Process

### v2

@@ 1 The Outlyingness Scores Based on Cluster Catch Digraphs
1.1 Introduction and Motivation
We have introduced several outlier detection methods based on RK-CCDs or UN-CCDs in the previous part, which share a common theme: (1) Clustering Formation: Identify potential clusters and construct a single or multiple covering balls around the center, covering majority points. (2) Outlier Identification: Apply the Density-based Mutual Catch Graph (D-MCG) method within each cluster to isolate points distant from the center and label them as outliers. Evaluation results demonstrated the effectiveness of these methods, with the SUN-MCCD method exhibiting superior overall performance. However, the outlyingness is not a binary property [breunig2000lof]. Moreover, the CCD-based methods do not provide a measure of outlyingness. An Outlyingness Score (OS) quantifies the degree of outlyingness for each observation. Incorporating OSs offers several advantages [kriegel2011interpreting]: (a) OSs enhance the interpretability of outlier detection results. (b) Ranking observations by their OSs enables user-defined thresholds for outlier classification. (c) OSs contribute to a richer understanding of the data distribution and point patterns. Some established, well-known OS-based outlier detection methods include the following: (i) Isolation Forest employs multiple decision trees to compute anomaly scores [liu2008isolation]; (ii) Local Information Graph-based Random Walk model (LIGRW) identifies outliers by analyzing unusual patterns in random walks [wang2018new]; (iii) Local Outlier Factor (LOF) is a prototype that utilizes local reachability density to compute outlyingness [breunig2000lof]; (iv) Connectivity-based Outlier Factor (COF) enhances LOF by addressing limitations in detecting outliers with similar densities but differing neighbor patterns [tang2002enhancing]; (v) LOcal Correlation Integral (LOCI) introduces a threshold-based score measuring the intensity of points within a given radius [papadimitriou2003loci]; (vi) Local Outlier Probabilities (LoOP) assigns an outlier probability based on the deviates of a point from its local context [kriegel2009loop]; (vii) Outlier Detection using In-degree Number (ODIN) identifies outliers as points with low in-degree numbers in a kkNN graph [hautamaki2004outlier]. Other OS-based methods include Angle-Based Outlier Detection (ABOD) [kriegel2008angle], Histogram-based Outlier Score (HBOS) [goldstein2012histogram], and Feature Bagging [lazarevic2005feature].
@@ 1.2 Outbound Outlyingness Score
Given a dataset $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}$ and a CCD constructed on it, the outbound neighbors of a point $x_{i}\in\mathrm{X}$, denoted as $N_{O}(x_{i})$, are those points that are covered by the covering ball centered at $x_{i}$, i.e.,
@@
Given a dataset $\mathrm{X}$ and a point $x_{i}\in\mathrm{X}$, the vicinity density of $x_{i}$, denoted by $\rho_{x_{i}}$, is defined as follows,
$\rho_{x_{i}}:=\frac{|\{x_{j}|x_{j}\in B(x_{i},r_{x_{i}})\}|}{r_{x_{i}}}^{\frac{1}{d}},$ (2)
@@
Given a dataset $\mathrm{X}$ and a point $x_{i}\in\mathrm{X}$, an Outbound Outlyingness Score of $x_{i}$, denoted as $OOS(x_{i})$, is given as follows,
@@ 1.2.1 An illustration of OOS with UN-CCDs
In Figure 1(b), we construct UN-CCD to illustrate the computation of OOS. While OOS is independent of the clustering result, it is worth noting that the UN-CCD method assigns outliers $O_{1},O_{2},...,O_{6}$ to $C_{1}$, $O_{7}$ and $O_{8}$ to $C_{3}$, and $O_{9}$ to $C_{2}$. We first focus on the cluster $C_{1}$. Figure 1(b) presents the covering balls of $O_{5}$ and $O_{6}$, along with their outbound neighbors $P_{i}\ (i=1,2,3)$. $P_{i}^{\prime}s$ vicinity densities are substantially higher than those of $O_{5}$ and $O_{6}$, resulting in OOSs values of 2.47 and 10.14 for $O_{5}$ and $O_{6}$, respectively. In contrast, most of the remaining points within $C_{1}$ exhibit OOS values below 2 (Figure 1(d)). Consequently, $O_{5}$ and $O_{6}$ are clearly distinguished by their markedly higher OOS values within $C_{1}$. Similarly, $O_{9}$ receives a much higher OOS than the other points in $C_{2}$.
@@
However, as one of the limitations, OOS is not effective on collective outliers, showing little robustness to the masking problem.
@@ 1.3 Inbound Outlyingness Score
Note that the larger $CI(x_{i})$, the less $x_{i}$ deviates from other points (i.e., the more likely $x_{i}$ belongs to the cluster). So, we may define the Inbound Outlyingness Score (IOS) of $x_{i}$ as the reciprocal of $CI(x_{i})$. Furthermore, to avoid a zero denominator when $CI(x_{i})$ is 0 (i.e., when a point does not have any inbound neighbors), we add $\rho(x_{i})$ to $CI(x_{i})$, ranking the points without any inbound neighbors by the reciprocal of their vicinity densities.
@@ 1.3.1 An illustration of IOS with UN-CCDs
By the “three-sigma rule” [maronna2019robust], points with standardized scores larger than 3 can typically be deemed as outliers. However, this traditional measure is not robust to outliers, as the scores of outliers have a substantial and unbounded influence on the sample mean and the sample SD [maronna2019robust], which is particularly problematic in an outlier detection method. Fortunately, there are robust alternatives to mean and SD, and we employ median (Med) and the Normalized Median Absolute Deviation about the median (MADN). With this notion, a robust version of Equation (7) can be defined as follows.
@@
where $MADN(IOS(C))=Med\{IOS(C)-Med(IOS(C))\}/0.6745$. It is worth noting that the MADN of a random variable following $N(\mu,\delta)$ distribution is $\delta$ [maronna2019robust]. In the remaining sections, all IOS values will be standardized by default, i.e., $IOS(x_{i})=IOS^{std}(x_{i})$.
@@
Suppose the IOSs of a cluster $C$ are ranked in ascending order as follows:
@@
where the IOS of, WLOG, $x_{1}$, $x_{2}$, …, $x_{m}$ equal to $IOS_{(k)}$ ($m\leq n_{c}$). To break ties, we assign new scores to them in a linearized fashion using $\rho_{x_{1}}$, $\rho_{x_{2}}$, …, $\rho_{x_{m}}$. Without loss of generality, for any point $x_{i}$ (i=1,,mi=1,...,m), the standardized IOS of it, denoted as $\widetilde{IOS}(x_{i})$, is defined as follows,
@@
  • Effective on collective outliers and robust to the masking problem.
@@
2 Computational Complexity of OOS and IOS
Time Complexity:
After the initial CCDs have been constructed, both the OOS and IOS take an additional $O(n^{2})$ time at most, dominated by the initial step of constructing the underlying CCDs. We justify their additional time complexity as follows,
  • Vicinity Density: The first step is to calculate the vicinity density for each of the $n$ points. This requires iterating through the neighbor list of each point. This step takes at most $O(n^{2})$ time.
  • OOS Calculation: For each point, OOS finds its “outbound neighbors” to compute the average vicinity density, and compares to the vicinity density of that point. which is at most $O(n^{2})$ for the entire dataset.
  • IOS Calculation: Calculating IOS requires summing the vicinity densities of all “inbound neighbors” for each point. This involves identifying which points’ covering balls contain a given point. This process, along with the subsequent standardization, also has a time complexity of at most $O(n^{2})$.
Therefore, both OOS and IOS take an additional $O(n^{2})$ time at most.
Space Complexity:
Similarly, the space complexity is determined by the underlying construction process of the CCD. Both OOS and IOS require additional space for vicinity density and outlyingness scores, which take $O(n)$ space, negligible in comparison to the $O(n^{2})$ space taken by CCDs. Therefore, the space complexity of OOS and IOS is $O(n^{2})$.
3 Monte Carlo Experiments with Synthetic Datasets
3.1 General Settings
We evaluate the performance of all the OSs by conducting Monte Carlo experiments with diverse simulation settings. The simulation settings are elaborated in [shi2024outlier]. These settings involve different factors (e.g., dimensionality, dataset sizes, cluster volumes, etc.), which vary among datasets. Moreover, there are two major types of simulation settings, one with only uniform clusters and the other with only Gaussian clusters. The goal is to evaluate four newly proposed CCD-based OSs against four existing CCD-based methods [shi2024outlier], seeking evidence of performance improvements.
Preliminary simulations on artificial data identified “elbow” cutoff points, serving as the thresholds for outlier detection. These thresholds, specific to each OS and dimensionality, are presented in Tables 1 and 2. Notably, the datasets with Gaussian clusters require higher thresholds than those with uniform ones, because we want to differentiate outliers from regular data points around the tails of the Gaussian distributions.
@@ 3.1 General Settings
We select True Positive Rate (TPR), True Negative Rate (TNR), Balanced Accuracies (BA), and $F_{2}$-score as the performance metrics. TPR is the proportion of outliers detected; TNR is the proportion of regular points correctly labeled, BA is the mean of TPR and TNR, $F_{\beta}$-score is the weighted harmonic mean of recall (TPR) and precision, prioritizing positive observations (outliers) [sasaki2007truth]. Since (in our opinion) recall is more important than precision in outlier detection, we set $\beta$ to 2, making recall twice as important. We emphasize the latter two measures for enhanced accuracy since the size of outliers and regular points is highly imbalanced. Comprehensive simulation results are provided in Tables 3 to 6. For better visualization, we also provide the simulation results in four line plots (Figures 4 to 7), presenting the performance trend with varying dimensions and data sizes.
@@
Table 3: Summary of the TPRs and TNRs of all the CCD-based OSs, with the simulation settings (with uniform clusters) elaborated in [shi2024outlier].
Figure 4: The line plots of the TPRs and TNRs of all CCD-based OSs, under the simulation settings (with uniform clusters) elaborated in [shi2024outlier].
@@
Table 4: Summary of the BAs and $F_{2}$-scores of all the CCD-based OSs, with the simulation settings (with uniform clusters) elaborated in [shi2024outlier].
Figure 5: The line plots of the BAs and $F_{2}$-scores of all CCD-based OSs, under the simulation settings (with uniform clusters) elaborated in our previous work [shi2024outlier].
@@
We first look at RKCCD-OOS and UNCCD-OOS. They exhibit similar performance across different dimensions. When $d\leq 20$, both OOSs achieve high BAs and $F_{2}$-scores, often exceeding 0.95 and 0.90, respectively, when the size of the dataset ($n$) is less than or equals to 200. However, as $n$ exceeds 500, the masking problem arises due to the increasing number of outliers. This issue is particularly prominent in lower dimensions where outlier intensities are high. For instance, with $d=2$ and $n=50$, UNCCD-OOS attains a BA of 0.980 and an $F_{2}$-scores of 0.932, but they drop to 0.720 and 0.475, respectively, as $n$ reaches 1000. Conversely, in high-dimensional settings ($d\geq 50$), both OOSs achieve higher TPRs with larger datasets. For instance, when $d=100$, the TPR of RKCCD-OOS progressively increases from 0.535 to 0.994 as $n$ increases. This trend can be attributed to the fact that in high-dimensional spaces with limited data, the vicinity densities of regular observations and outliers may not differ substantially. Recall that OOS measures the relative vicinity density of a point with its outbound neighbors, the reduced performance with smaller datasets is expected. Furthermore, when $d=50$ and $n=50$, the TPR of UNCCD-OOS is 0.957, substantially higher than that of RKCCD-OOS (0.535). This is because the covering balls of UN-CCDs are typically larger than those of RK-CCD, resulting in considerable differences between the vicinity densities between regular observations and outliers, which ultimately enhances the performance of OOS.
@@
Second, we evaluate the performance of RKCCD-IOS and UNCCD-IOS. With $d\leq 20$, both IOSs demonstrate comparable performance to OOSs when the data size $n$ is small. Moreover, an improvement is observed with larger $n$ values when compared to the two OOSs. For instance, with $d=3$ and $n\leq 200$, UNCCD-IOS achieves $F_{2}$-scores comparable to those of UNCCD-OOS (0.903, 0.914, and 0.904 versus 0.954, 0.947, and 0.922). However, when $n\geq 500$, UNCCD-IOS shows substantially higher $F_{2}$-scores than those of UNCCD-OOS (0.887, 0.848 versus 0.837, 0.709). This stems from the inherent advantage of IOS, which measures the cumulative influence a point receives from other points, and the cumulative influence of outliers is typically low compared to regular observations, making IOS more robust against the masking problem, particularly in simulation settings with many outliers. For example, with $d=2$ and $n=1000$, the $F_{2}$-scores of UNCCD-OOS and UNCCD-IOS are 0.475 and 0.833, respectively, showing a huge difference. However, RKCCD-IOS does not perform well when $d\geq 50$. For example, when $d=50$ and $n=1000$, the $F_{2}$-score of RKCCD-IOS is merely 0.605, much lower than that of RKCCD-OOS, which is 0.847. A reasonable explanation is the inherent challenges of high-dimensional clustering. As $d$ increases, the average distances between outliers and clusters grow, and the covering balls constructed by RK-CCDs are not large enough for high dimensions [shi2024outlier]. Consequently, as the number of outliers increases, RK-CCDs may falsely identify a set of close outliers as a valid cluster since they are not robust to the masking problem. Furthermore, IOS identifies an outlier by within-cluster comparisons, which does not mitigate this problem.
@@
Next, we compare the simulation results of the CCD-based OSs with previously proposed CCD-based outlier detection methods (which can be found in our previous work [shi2024outlier]).
@@
Following the analysis of simulations with uniform clusters, we explore the results obtained from simulations with Gaussian clusters (see Tables 5 and 6). Similarly, we begin by assessing the performance of the four OSs, and then compare them with the previously proposed CCD-based outlier detection methods (see our previous work [shi2024outlier]).
@@
Table 5: Summary of the TPRs and TNRs of all the CCD-based OSs, with the simulation settings (with Gaussian clusters) elaborated in our previous work [shi2024outlier].
Figure 6: The line plots of the TPRs and TNRs of all CCD-based OSs, under the simulation settings (with Gaussian clusters) elaborated in our previous work [shi2024outlier].
@@
Table 6: Summary of the BAs and $F_{2}$-scores of all the CCD-based OSs, with the simulation settings (with Gaussian clusters) elaborated in our previous work [shi2024outlier].
Figure 7: The line plots of the BAs and $F_{2}$-scores of all CCD-based OSs, under the simulation settings (with Gaussian clusters) elaborated in our previous work [shi2024outlier].
@@
Next, we consider RKCCD-IOS and UNCCD-IOS. Both IOSs achieve high $F_{2}$-scores when $d\leq 20$, because both IOSs address the masking problem raised by high-intensity outliers well. However, the performance of RKCCD-IOS drops substantially when $d\geq 50$, particularly when $n\geq 500$, For instance, when $d=100$ and $n=1000$, the $F_{2}$-score of RKCCD-IOS is 0.521. This limitation stems from the same issue encountered when using RK-CCDs for clustering, as detailed in [shi2024outlier].
Then, we compare the simulation results of the OSs with the U-MCCD and UN-MCCD methods. (whose performance can be found in our previous work [shi2024outlier])
@@
Next, we compare UNCCD-OOS, UNCCD-IOS, and the UN-MCCD method, which show varying performance. The UN-MCCD method, while superior to U-MCCD, exhibits limitations because it constructs a single covering ball for each cluster, which is inadequate to cover a Gaussian cluster. In contrast, both UNCCD-OOS and UNCCD-IOS demonstrate substantial improvement. Although the $F_{2}$-score of UNCCD-OOS may be as low as 0.293 (e.g., when $d=2$ and $n=1000$) due to the masking problem under lower dimensions, and occasionally underperforms compared to the UN-MCCD method, UNCCD-IOS consistently outperforms the UN-MCCD method and achieves considerably higher $F_{2}$-scores across all the simulation settings.
@@
Based on our simulation results, we observe that RKCCD-OOS and UNCCD-OOS deliver good performance except when $d\leq 5$ and $n$ is large, where the masking problem emerges. In contrast, RKCCD-IOS and UNCCD-IOS perform much better with lower dimensions but may degrade when $d$ is large because both RK-CCDs and UN-CCDs (especially RK-CCDs) may identify a set of close outliers as valid clusters in high dimensions. Compared to the U-MCCD and UN-MCCD methods, the CCD-based OSs return fewer false positives, achieving much higher $F_{2}$-scores in general. Moreover, they show substantial advantages in higher-dimensional datasets or datasets with Gaussian clusters, making them better choices in those cases.
3.2 Under Random Cluster Process